OpenStudy (anonymous):

How woul I express F(X)=x+2 and G(X)=3x+5 if the set up is (F times G)(x)?

OpenStudy (anonymous):

F(X)=x+2 and G(X)=3x+5 if F times G)(x)? (x+2)(3x+5)=3x^2 +6x+5x+10 =3x^2 +11x+10